Sheridan Monograms sells stadium blankets that have been monogrammed with high school and university emblems. The blankets retail for $46 throughout the country to loyal alumni of over 3,800 schools. Sheridan’s variable costs are 42% of sales; fixed costs are $116,000 per month.

Assume that variable costs increase to 47% of the current sales price and fixed costs increase by $11,900 per month. If Sheridan were to raise its sales price by 12% to cover these new costs, what would be the new annual breakeven point in sales dollars?
Answer #1

Solution:

New Variable costs per unit = $46 * 47% = $21.62

New Selling Price = $46*112% = $51.52

Contribution margin per unit = (51.52 - 21.62) / $51.52 = 58.04%

breakeven point in sales dollars = ($116000+$11900) / 58.04% = $220,365